A few years ago, an analysis of some Wikipedia reference pages noted that the references (citations) at the ends of articles are tend to break over time. That is, the links to the cited articles or sites can’t be followed. It was claimed that 3.9% of these references are incorrect within a span of about one year. If a particular Wikipedia article contains seven references, what is the probability that all seven references are still valid one year later? What assumption must you make in order to calculate this probability?
1
Expert's answer
2015-02-25T10:14:58-0500
If it was claimed that 3.9% of the references are incorrect within a span of about one year, and if a particular Wikipedia article contains seven references, then by multiplication rule of probability, the probability that all seven references are still valid one year later is given byP = (1 - p)^7 = (1 - 0.039)^7 = 0.961^7 = 0.7569 or 75.69 %.In order to calculate this probability, we should make the next assumptions:
1) probability for every article to be valid is the same;2) an incorrect reference does not affect the probability that other article is incorrect.